wat is 29-58? inoe its -29 but i want to know the steps on how to get to get -29

Answer by drj(1380)

What is 29-58? I know it's -29 but I want to know the steps on how to get to get -29.


29-58=29+(-58)=-29. Where we used the additive inverse of 58.

Think of walking 29 forward steps and then 58 backward steps. The net result is that you went 29 steps backwards from where you started.

I hope the above steps were helpful.

For free Step-By-Step Videos on Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra or for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J